Circular No. 6538

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997E IN NGC 2258 S. Nakano, Sumoto, Japan, reports the discovery by Reiki Kushida,

Yatsugatake South Base Observatory, of an apparent supernova (mag 15.7) on seven unfiltered CCD frames taken around Jan. 14.530 UT with a 0.40m f/5 reflector. Y. Kushida provides the following measurement of the position of SN 1997E from the discovery frames: α = 6^{h}47^{m}38^{s}.10, δ = +74^{o}29′51′′.0 (equinox 2000.0), which is 32′′ west and 57′′ north of the center of NGC 2258. There is no star at the position of the new star on Kushida’s frames taken on 1996 Dec. 15, 29, and 1997 Jan. 4, nor is anything present on the Digital Sky Survey.

P. Garnavich and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, communicate: “Spectrograms of SN 1997E were obtained by J. Peters on Jan. 16.2 UT at the F.L. Whipple Observatory (FLWO) 1.5-m Tillinghast telescope, showing this to be a type-Ia supernova at maximum. The relative strengths of the Si II features at 617.9 and 582.2 nm, and the narrow Si II feature at 491.3 nm, imply that this supernova may have a fastdeclining light curve similar to that of SN 1992A. CCD images obtained on Jan. 16.4 by J. Raymond with the FLWO 1.2-m telescope yield visual mag 15.6 for SN 1997E.

NEW X-RAY BURSTER L. Piro, Istituto Astrofisica Spaziale (IAS), CNR, Frascati; J. Heise

and R. Jager, Space Research Organization of the Netherlands (SRON), Utrecht; M. Feroci, IAS; G. D’Andreta, G. Spoliti, and A. Coletta, BeppoSAX, Rome; H. Muller, BeppoSAX, Rome, and SRON; and D. Ricci, BeppoSAX, Rome, write: “On Jan. 13.295011 UT, the BeppoSAX Wide Field Camera No. 1 (WFC1) detected an x-ray burst in the range 2–20 keV, seen in the ratemeter as a single peak and showing a fast rise (on the order of 10 s) with an exponential decay lasting ∼ 150 s. This event is likely to be an x-ray burster and not a γ-ray burst, as it was not seen by the BeppoSAX Gamma Ray Burst Monitor. The derived position (α = 12^{h}49^{m}50^{s}, δ = −59^{o}10′.8, equinox 2000.0; estimated 10′ error radius) does not correspond to any catalogued source. The peak flux corresponds to ∼ 1.9 Crab. The WFC1 observed the same field for ∼ 40 000 s before and 2500 s after the event, with no positive detection, leading to an upper limit to the steady flux from the burst source of ∼ 10 mCrab.”

1997 January 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green